OpenStudy (anonymous):

Find the supplementary angle of 34°. (Points : 5) 146° 56° 34° 68°

jimthompson5910 (jim_thompson5910):

Supplementary angles add to 180 degrees. So this means that 34+x= 180 and x = 180-34 = 146 So the answer is choice A